# A Survey of Generative AI for Intelligent Transportation Systems

### v1

@@ 5.2. Traffic Scenario Generation
Vehicle trajectory generation. Many researchers use generative AI technology to generate more vehicle trajectory data (Krajewski et al., 2018, 2019; Choi et al., 2021; Liao et al., 2022; Zhang et al., 2022e; Chen et al., 2021c; Ozturk et al., 2020; Kumar et al., 2023; Sun et al., 2023; Xu et al., 2023d). In (Krajewski et al., 2018), the authors introduced trajectory GAN and trajectory VAE to generate realistic synthetic lane change maneuver trajectories without expert knowledge. Diff-Traj (Zhu et al., 2023) was proposed for generating GPS trajectories by effectively learning spatio-temporal features from historical trajectories and leveraging the generation capability of the diffusion model. Xu et al. (Xu et al., 2023b) designed a generative AI-based autonomous driving architecture that utilizes large text-to-image models based on DPM to generate conditioned traffic and driving data in simulations. Differently, some searchers focused on the controlled vehicle trajectory generation (Arnelid et al., 2019; Zhong et al., 2023a, b). For example, the authors in (Arnelid et al., 2019) introduced a recurrent conditional GAN model that generates sensor errors with long-term temporal correlation in both the generator and the discriminator networks. Zhong et al. (Zhong et al., 2023a) introduced a scene-level conditional diffusion model that incorporates a spatio-temporal transformer architecture to produce both realistic and controllable traffic data.
Human trajectory generation. Regarding human trajectory generation, Xiong et al. (Xiong et al., 2023) proposed a semantic-guiding GAN model equipped with semantic knowledge to generate human trajectories. Rempe et al. (Rempe et al., 2023) studied controlled pedestrian trajectory generation. They utilized guided diffusion modeling for test-time controllability, and integrating it with a physics-based humanoid controller for a closed-loop system that places large crowds in various terrains while meeting user-defined goals.
@@
Anomaly data generation. Anomaly data such as traffic accident and crash data is essential in autonomous driving for safety assessment and scenario testing. However, this type of traffic data is hard to collect. For example, as illustrated in (Cai et al., 2020), the ratio of crash data to non-crash data is 1 to 11,000, indicating a significant imbalance. Several works use generative AI techniques to address this problem (Islam et al., 2021; Chen et al., 2021e, 2022a; Huo et al., 2021; Zarei and Hellinga, 2021). For instance, Islam et al. (Islam et al., 2021), employed VAE to encode events into a latent space, from which they subsequently sampled to generate crash data. In (Huo et al., 2021), a text-to-traffic GAN model was proposed, which integrates traffic data with semantic information collected from social media to produce traffic situations. Zarei et al. (Zarei and Hellinga, 2021) developed a conditional GAN-based model that utilizes crash count as a condition to generate crash data.
@@ 5.3. Traffic Flow Generation
Traffic flow generation refers to generate vehicular traffic volumes or crowd flow within a specific area or between regions in a transportation network. Numerous works adopt generative AI techniques to address this task (Chen et al., 2021a; Li et al., 2022d; Zhou et al., 2023; Rong et al., 2023a, b). Specifically, Li et al. (Li et al., 2022d) presented an attentive dual-head spatial-temporal GAN model for generating crowd flow. This model incorporated an attentive mechanism for tracking temporal changes and a self-attention network to capture intricate spatiotemporal crowd flow patterns, and it adopted a dual-head discriminator that implements a training strategy with two distinct objectives to mitigate the adverse effects of rapid overfitting. Zhou et al. (Zhou et al., 2023) constructed an urban knowledge graph to extract region features and model the urban environment, and they subsequently designed a knowledge-aware spatio-temporal diffusion model for generating urban flow. In (Rong et al., 2023a), the authors employed a graph denoising diffusion technique to generate urban OD flow, enabling the learning of how the likelihood of nodes and edges are interconnected in the OD network.
@@ 6. Generative AI for Traffic Decision-making
In autonomous driving scenarios, traffic decision-making is a series of decision-making processes made by vehicles based on current dynamic traffic environment and task requirements when driving on the road, such as driving routes, speeds, driving modes and so on. It is a challenging task for ensuring that vehicles make safe and accurate decisions in complex environments. The first challenge we face is dealing with the dynamic and unpredictable nature of traffic environments. These environments are constantly changing, with various factors like road conditions, unexpected moves by human drivers, and sudden decisions by pedestrians and cyclists. For example, a driver might suddenly brake or change lanes without warning, or pedestrians might cross the road unexpectedly. These uncertainties compel autonomous systems to be perpetually alert and adaptive, which becomes computationally intensive and requires robust prediction algorithms. The second challenge involves rare but critical events that can test the safety of autonomous vehicles, such as a child running onto the road or a mechanical failure causing loss of control. Gathering real-world data on these rare events is difficult but crucial for creating a comprehensive training dataset. Without exposure to these scenarios, autonomous systems may not know how to handle them properly, potentially leading to safety risks. The third challenge relates to the limitations of sensors like LiDAR, RADAR, and cameras. These sensors have restrictions in terms of their range, resolution, and susceptibility to environmental factors like fog, rain, and snow. These limitations can sometimes lead to uncertainty or incorrect decision-making by the autonomous vehicle, which can compromise both safety and efficiency.
@@
Traditional decision-making methods, such as rule-based methods (McNaughton et al., 2011; De Beaucorps et al., 2017) and game theory-based methods (Wang et al., 2015; Hang et al., 2020; Zhao et al., 2021a), often have limited effectiveness in dealing with complex traffic situations, while reinforcement learning (RL), as a flexible, efficient, and powerful method, has gained ground in autonomous driving decision-making. However, its general application is hindered by training complexity, especially in handling multiple tasks. Generative AI technology holds promise in tackling these challenges. Leveraging their robust generative capabilities, they can generate realistic training scenarios mimicking the unpredictability of real-world driving, reducing the need for extensive road testing. Moreover, they can produce rare traffic events to ensure autonomous systems are robust against critical scenarios. Additionally, generative AI can enhance sensor data, helping vehicles adapt to challenging environmental conditions.
Recently, researchers have shifted their focus to using generative AI techniques in traffic decison-making. Wang et al. (Wang et al., 2022) proposed an autonomous decision-making approach for high-speed trains, integrating GAN-based learning approach and a distributed tracking controller. Addressing data insufficiency, they introduced a GAN-based data augmentation method to generate time series samples for training purposes. They then developed a trajectory prediction network that compute the reference speed trajectory in real-time conditions. Further, they designed a distributed tracking control model, utilizing model predictive control and dual decomposition, to optimize train movement.
Liu et al. (Liu et al., 2023a) proposed a multi-task decision-making GPT model for autonomous driving at unsignalized intersections. By leveraging the exceptional learning capabilities and computational efficiency of RL, this model had ability to simultaneously manages multiple driving tasks. The authors first trained RL expert models for single decision-making tasks, and then used the sampled expert data to guide the offline GPT training. As a result, the model outperforms conventional single-task RL decision-making models.

### v2

@@ 5.2. Traffic Scenario Generation
Vehicle trajectory generation. Many researchers use generative AI technology to generate more vehicle trajectory data (Krajewski et al., 2018, 2019; Choi et al., 2021; Liao et al., 2022; Zhang et al., 2022e; Chen et al., 2021c; Ozturk et al., 2020; Kumar et al., 2023; Sun et al., 2023; Xu et al., 2023d). In (Krajewski et al., 2018), the authors introduced trajectory GAN and trajectory VAE to generate realistic synthetic lane-change maneuver trajectories without expert knowledge. This approach not only facilitated the creation of diverse training datasets for autonomous driving systems but also reduced the dependency on extensive expert annotations. PS-TrajGAIL (Wang et al., 2024a) modeled urban vehicle trajectory generation as a partially observable Markov decision process within a GAIL-based structure. This approach leveraged IL to capture the complex decision-making processes of human drivers, thereby generating synthetic trajectories that exhibited high fidelity to real-world behaviors. Experimental evaluations on both synthetic and real-world datasets demonstrated that PS-TrajGAIL significantly outperforms existing baselines in terms of trajectory realism and predictive accuracy. Diff-Traj (Zhu et al., 2023) was designed for generating GPS trajectories by learning spatio-temporal features from historical trajectories. By employing DPMs, this model captured the stochastic and dynamic nature of human driving behaviors, producing high-quality synthetic trajectories that retain the statistical properties of real-world data. WcDT (Yang et al., 2024c) combined DPMs with transformers to generate diverse and realistic autonomous driving trajectories. By encoding historical data and environmental features through transformer-based encoders, WcDT effectively captured multifaceted interactions and temporal dependencies present in driving scenarios. The integration of diffusion models enhanced the framework’s ability to produce high-fidelity trajectories that reflect the variability and unpredictability of human driving behavior.
@@
Differently, some searchers focused on the controlled vehicle trajectory generation (Arnelid et al., 2019; Zhong et al., 2023a, b). For example, the authors in (Arnelid et al., 2019) proposed RC-GAN, a recurrent conditional GAN model that incorporated LSTM in both the generator and discriminator, enabling the generation of sensor errors with long-term temporal correlations. Empirical evaluation on real road data demonstrated that RC-GAN produced significantly more realistic sensor error simulations. Zhong et al. (Zhong et al., 2023a) introduced a scene-level conditional diffusion model that incorporated a spatio-temporal Transformer architecture to produce realistic and controllable traffic data. A key innovation was incorporating language instructions through a LLM, which translated user queries into loss functions that guide the diffusion model. Evaluations demonstrated that this approach not only produced realistic traffic patterns but also effectively adhered to user constraints, outperforming existing methods in both realism and controllability. Xu et al. (Xu et al., 2023b) designed a generative AI-based autonomous driving architecture that utilizes large text-to-image models based on DPM to generate conditioned traffic and driving data in simulations. The approach improved simulation fidelity, providing a robust tool for testing and validating autonomous driving systems.
Human trajectory generation. Regarding human trajectory generation, Xiong et al. (Xiong et al., 2023) proposed a semantic-guiding GAN model that integrates semantic knowledge to generate realistic human trajectories. By employing an attention-based generator and a CNN-based discriminator, this approach effectively captured both the sequential and spatial dependencies of human movements. Experiments validated its utility in simulating epidemic spread scenarios, achieving high correlation with real data. STAGE (Cao et al., 2024) employed a multi-task GAN with spatiotemporal knowledge, integrating a multi-stage generator and spatial consistency loss to produce privacy-preserving synthetic trajectory. Experimental results indicated that STAGE outperformed baseline models in data distribution alignment, making it a robust tool for applications requiring high data fidelity and privacy protection. G-GAIL (Li et al., 2024b) integrated game theory with GAIL to model pedestrian-vehicle interactions as a Markov decision process, enabling realistic simulation of pedestrian behaviors and enhancing autonomous driving scenario testing. The model demonstrated superior performance in replicating pedestrian behaviors under various conditions, enhancing the reliability of autonomous driving scenario testing. Rempe et al. (Rempe et al., 2023) introduced a controlled pedestrian trajectory generation approach using guided diffusion modeling and a physics-based humanoid controller, generating realistic movements based on user-defined goals and environmental constraints, thus improving the realism and flexibility of simulations.
@@
Anomaly data generation. Anomaly data such as traffic accident and crash data is essential in autonomous driving for safety assessment and scenario testing. However, this type of traffic data is hard to collect. For example, as illustrated in (Cai et al., 2020), the ratio of crash data to non-crash data is 1 to 11,000, indicating a significant imbalance. Several works use generative AI techniques to address this problem (Islam et al., 2021; Chen et al., 2021e, 2022a; Huo et al., 2021; Zarei and Hellinga, 2021). For instance, Islam et al. (Islam et al., 2021) employed VAE to encode imbalanced crash and non-crash events into a latent space for generating realistic crash data. This approach not only achieved statistical similarity with real crash data but also significantly improved prediction model performance. In (Huo et al., 2021), a text-to-traffic GAN model was proposed, which integrates traffic data with semantic information collected from social media. By incorporating a global-local loss to bridge the modality gap between textual and traffic data, their model achieved superior authenticity in generated traffic scenarios. Zarei et al. (Zarei and Hellinga, 2021) developed a conditional GAN-based model that utilizes crash counts as conditions to generate realistic crash data. Their method demonstrated superior performance in hotspot identification, prediction accuracy, and dispersion parameter estimation, especially in low-dispersion scenarios.
@@ 5.3. Traffic Flow Generation
Traffic flow generation refers to generate vehicular traffic volumes or crowd flow within a specific area or between regions in a transportation network. Numerous works adopt generative AI techniques to address this task (Chen et al., 2021a; Li et al., 2022d; Zhou et al., 2023; Rong et al., 2023a, b). Specifically, Li et al. (Li et al., 2022d) presented an attentive dual-head spatial-temporal GAN model for generating crowd flow data. This model incorporated an attentive mechanism for tracking temporal changes and a self-attention network to capture intricate spatiotemporal crowd flow patterns, and it adopted a dual-head discriminator that implements a training strategy with two distinct objectives to mitigate the adverse effects of rapid overfitting. Zhou et al. (Zhou et al., 2023) developed a knowledge-aware spatio-temporal diffusion model that utilizes an urban knowledge graph to generate dynamic urban flow data. By integrating region features and environmental factors, this model accurately generated urban flows for regions without historical data. In (Rong et al., 2023a), the authors employed a graph denoising diffusion technique to generate urban OD flow, effectively capturing the interconnections of nodes and edges within large urban networks. By decomposing the generative process into network topology and edge weight generation, their model achieved superior realism in OD matrices across multiple cities.
5.4. Discussion
The literature review highlights several key lessons regarding the application of generative AI techniques in traffic simulation. Firstly, generative models in (Zarei and Hellinga, 2021; Islam et al., 2021) effectively address data imbalance by synthesizing realistic crash and rare event data. Secondly, integrating multi-modal and semantic information, as seen in $\text{T}^{2}$ (Huo et al., 2021), enriches traffic scenario generation by combining traffic data with social media insights, resulting in more authentic and contextually relevant simulations. Moreover, privacy-preserving techniques like STAGE (Cao et al., 2024) ensure data fidelity while protecting individual privacy, making synthetic data suitable for sensitive applications. However, these advanced models often face challenges related to computational complexity and scalability, highlighting the need for further optimization to enable real-time and large-scale implementations.
@@ 6. Generative AI for Traffic Decision-making
Traffic decision-making involves a series of decision-making processes made by vehicles or other road users based on dynamic traffic environment and specific task requirements. Among these processes, route planning, traffic signal control (TSC) and vehicle control are three fundamental aspects. Route planning generally refers to generating an optimal path from a designated starting point to a destination, considering factors such as departure time, road conditions, and traffic patterns. TSC plays a crucial role in coordinating the movement of vehicles and pedestrians at intersections, thereby reducing traffic congestion. Vehicle control involves managing the vehicle’s operation, including actions such as acceleration, steering, and braking, to navigate safely along driving route.
Figure 7. The relation between different topics in generative AI for traffic decision-making research.
@@
Ensuring that route planning, TSC, and vehicle control collaboratively make safe and accurate decisions in complex environments is a formidable challenge. One significant challenge in traffic decision-making lies in the unpredictable dynamics of traffic environments. These environments are constantly changing due to various factors, including road conditions, unexpected moves by human drivers, and sudden decisions by pedestrians and cyclists. For example, a driver might suddenly brake or change lanes without warning, or pedestrians might cross the road unexpectedly. These uncertainties compel autonomous systems to be perpetually alert and adaptive, which becomes computationally intensive and requires robust prediction algorithms. Another critical challenge involves rare but critical events that can pose significant risks to safety, such as a child running onto the road or a mechanical failure causing loss of control. Collecting real-world data on these rare events is difficult but crucial for creating a comprehensive training dataset. Without exposure to these scenarios, autonomous systems may not know how to handle them properly, potentially leading to safety risks. Additionally, the limitations of sensors (e.g, LiDAR, RADAR, and cameras) present another obstacle. These sensors often have restrictions regarding their operational range and resolution, as well as their susceptibility to adverse weather conditions, including fog, rain, and snow. Such limitations can introduce significant uncertainties, resulting in suboptimal decision-making and compromising both safety and efficiency within traffic operations.
@@
Traditional decision-making methods, such as rule-based methods (McNaughton et al., 2011; De Beaucorps et al., 2017) and game theory-based methods (Wang et al., 2015; Hang et al., 2020; Zhao et al., 2021), often have limited effectiveness in dealing with complex traffic situations, while reinforcement learning (RL), as a flexible, efficient, and powerful method, has gained ground in traffic decision-making. However, the training complexity associated with handling multiple tasks limits its widespread applicability. Generative AI technology offers promising solutions to these challenges by leveraging its strengths in data generation, reasoning, and sequence modeling. For instance, its sequence modeling capabilities help in understanding and predicting complex driving patterns over time. Furthermore, generative AI’s strong analytical and reasoning abilities allow for more effective interpretation of external traffic conditions and instructions, thereby providing important guidance for accurate decision-making. To mitigate the impact of rare events, generative AI can generate realistic training scenarios mimicking the unpredictability of real-world driving, reducing the need for extensive road testing. By producing rare and critical traffic events, generative AI ensures autonomous systems are robust against challenging situations. Furthermore, generative AI can enhance sensor data, enabling vehicles to handle challenging environmental conditions with greater adaptability and accuracy.
In the following subsections, we conduct a comprehensive literature review and explore how generative AI can enhance decision-making processes within traffic systems, specifically focusing on route planning, traffic signal control and vehicle control. Their relationship is illustrated in Figure 7.
6.1. Route Planning
Generative approaches have become integral to route planning tasks within intelligent transportation systems, leveraging their robust data generation and powerful learning capabilities to incorporate global geographical knowledge.
One category of approaches utilizes GANs to produce realistic and efficient routes (Fu and Lee, 2021; Choi et al., 2022; Wang et al., 2022). For example, ProgRPGAN (Fu and Lee, 2021) employed a progressive route planning GAN to create realistic paths across multi-level map resolutions. Similarly, Choi et al. (Choi et al., 2022) developed a GAN-based framework that integrates feature extraction and path generation networks to produce multiple plausible routes from egocentric images without relying on high-definition maps.
Another category of approaches integrates RL with generative models to achieve expert-level path planning. MixGAIL (Yang et al., 2024a) combined transition-aware adversarial imitation learning and GAIL techniques within a mixed policy gradient actor-critic framework. Building on this, a VAE-GAN framework for path optimization (Wang, 2022) combined VAEs with GANs to address the vehicle routing problem on graphs. This approach utilized GNNs to model the complex relationships within road networks and employed variational reasoning to identify optimal nodes by dividing the graph into subgraphs centered around a root node. By integrating RL, the framework achieved end-to-end optimization, enhancing both learning efficiency and generalization capabilities.
In addition, the integration of LLMs represents a great advancement in providing explainable and adaptive path planning. DynamicRouteGPT (Zhou et al., 2024) utilized causal inference and pretrained language models to offer explainable and adaptive real-time path planning in complex traffic scenarios. Addressing the limitations of traditional algorithms in large-scale environments, LLM-A* (Meng et al., 2024) proposed a novel path planning method that synergistically combines the precise pathfinding capabilities of A* with the global reasoning abilities of LLMs. This hybrid approach aims to enhance pathfinding efficiency in terms of time and space complexity while maintaining path validity.
6.2. Traffic Signal Control
Generative methods greatly improve the effectiveness of TSC by improving data integrity, scalability, and decision-making efficiency. The existing literature on generative methods for TSC can be divided into two categories: single-intersection and multi-intersection control approaches.
Single-intersection control. Generative approaches improve single-intersection traffic signals by recovering missing data. For instance, DiffLight (Chen et al., 2024a) addressed missing data in TSC by combining data imputation with decision-making, using a diffusion model to ensure robust control even with incomplete data. Other research focuses focus on the integration of LLMs with RL (Wang et al., 2024b; Pang et al., 2024). The authors in (Wang et al., 2024b) introduced a hybrid framework that integrates LLMs with perception and decision-making tools, enabling the system to adapt to diverse traffic conditions without additional training. Pang et al. (Pang et al., 2024) further advanced this integration by employing LLMs to evaluate and adjust RL decisions under degraded communication and rare events. These integrative methods demonstrate substantial advantages in adaptability and resilience, effectively addressing the limitations inherent in traditional RL-based systems when confronted with dynamic and unpredictable traffic environments.
Multi-intersection coordination. This category focuses on data recovery and the coordination of traffic signals across multiple intersections. A decentralized adaptive TSC framework (Wang et al., 2021b) employed a GAN-based algorithm for efficient traffic data recovery. This approach significantly reduced vehicle travel time and maintained high throughput by enabling collaboration with neighboring intersections without heavy data exchange. Similarly, Xu et al. (Xu et al., 2024a) leveraged a Wasserstein GAN to estimate agent state spaces for multi-intersection settings, which effectively handles missing data and enhances system robustness. When integrated with a graph-based deep RL model, this approach outperformed baseline models in handling complex interactions and incomplete data in multi-intersection environments.
6.3. Vehicle Control
Vehicle control primarily involves exploring decision-making approaches in autonomous driving scenarios, including individual vehicle control and connected vehicle control strategies.
Individual vehicle. This subsection discusses methods to enhance decision-making for individual vehicles using LLM-assisted approaches and other generative techniques.
LLM-assisted approaches. The integration of LLMs into autonomous vehicle systems has emerged as a transformative approach to enhancing decision-making capabilities in data analysis, reasoning, and sequential decision modeling. By leveraging LLMs, these systems can effectively interpret and process multi-modal external data, perform complex reasoning akin to human commonsense, and generate coherent decision sequences that enhance navigation and safety. For instance, DriveGPT4 (Xu et al., 2024b) and TrafficGPT (Zhang et al., 2024) illustrated how LLMs can process multi-frame video inputs and traffic data to produce vehicle controls and decompose complex tasks, respectively. This capability significantly enhances the system’s ability to manage intricate driving scenarios. Additionally, frameworks proposed in works (Cui et al., 2024b) and (Cui et al., 2024a) incorporated environmental information and driver commands, enabling human-like reasoning and real-time optimization of driving maneuvers. These frameworks also personalized user experiences and foster trust through explainable decisions. Furthermore, cognitive reasoning pathways and translation algorithms, as employed in the literature (Sha et al., 2023), ensured seamless integration of high-level LLM decisions with low-level vehicle controls, thereby enhancing adaptability and reliability. Moreover, models such as the multi-task decision-making GPT (Liu et al., 2023a) and the sequence modeling approach within a constrained Markov decision process (Liu et al., 2024a) explored the capability of LLMs to handle multiple driving tasks simultaneously and improve exploration and safety through advanced techniques like entropy regularization. The use of graph-of-thought structures in explainable decision-making (Chi et al., 2024) highlighted the importance of transparency and trust by generating natural language rationales for driving choices. In summary, this integration not only addresses critical challenges such as safety assurance, generalization to rare events, and interpretability but also significantly improves the safety, efficiency, and reliability of autonomous driving, paving the way for more intelligent and trustworthy transportation solutions.
Other generative techniques. Besides LLM-aided approaches, other generative techniques provide great advancements in autonomous vehicle decision-making by enhancing robustness and safety in complex driving scenarios (Couto and Antonelo, 2024; Liu et al., 2024b; Yang et al., 2024b). For example, the authors in (Couto and Antonelo, 2024) propose a multi-level GAIL framework that uses a GAN to generate Bird’s-Eye View representations and a GAIL module to learn vehicle control from these representations, effectively decoupling representation learning from the driving task to achieve robust autonomous navigation. DDM-Lag (Liu et al., 2024b) used a diffusion-based generative process to model autonomous vehicle decision-making, incorporated Lagrangian-based safety enhancements, and integrated hybrid policy updates with safety-constrained optimization to enhance decision performance in complex driving scenarios. This approach addresses the critical need for safety assurance in AVs by mitigating risks associated with decision-making under uncertainty, ensuring reliable performance in intricate driving situations.
Connected vehicles. Recent advancements in connected vehicle control demonstrate the significant benefits of integrating LLMs into autonomous driving frameworks. CoMAL (Yao et al., 2024) leveraged LLMs to facilitate collaboration among autonomous vehicles, enhancing traffic flow and reducing congestion. A key component of CoMAL is the reason engine, which processes scenario descriptions and system prompts as inputs. Using a hierarchical chain-of-thought prompt, the LLM generated driving strategies based on the intelligent driver model, allowing each vehicle to respond effectively to dynamic situations. This approach not only outperformed traditional reinforcement learning methods on the Flow benchmark but also highlighted the strong cooperative capabilities of LLM-driven agents. Similarly, CoDrivingLLM (Fang et al., 2024b) employed LLMs within a reasoning module for decision-making, incorporating state perception, intent sharing, negotiation, and chain-of-thought processing to produce final actions. Experimental results showed that CoDrivingLLM surpasses rule-based, optimization-based, and machine learning methods, achieving higher success rates, effective conflict resolution, and balanced efficiency in complex multi-vehicle traffic scenarios.
6.4. Discussion
While existing literature has made significant progress in using generative AI to enhance the effectiveness of traffic decision-making, there are still areas for improvement in capturing the unpredictable dynamics of traffic environments, addressing rare but critical events, and overcoming the limitations of sensors. For instance, current LLM-aided methods such as DriveGPT4 (Xu et al., 2024b) and TrafficGPT (Zhang et al., 2024), although capable of processing multi-frame video and traffic data, still require enhancements in decision robustness and real-time responsiveness in unknown scenarios. Additionally, while methods like DiffusionES (Yang et al., 2024b) attempt to augment the flexibility and diversity of autonomous vehicle decision-making strategies, the diversity and authenticity of generating extremely rare events remain to be validated. Furthermore, the literature provides limited discussion on enhancing perception and decision-making capabilities under adverse environmental conditions. Addressing it will be a crucial area for future research.